# Universally Robust Control of Open Quantum Systems

### v1

@@ Abstract
Mitigating noise-induced decoherence is the central challenge in controlling open quantum systems. While existing robust protocols often require precise noise models, we introduce a universal framework for noise-agnostic quantum control that achieves high-fidelity operations without prior environmental noise characterization. This framework capitalizes on the dynamical modification of the system-environment coupling through control drives, an effect rigorously encoded in the dynamical equation. Since the derived noise sensitivity metric remains independent of the coupling details between the system and the environment, our protocol demonstrates provable robustness against arbitrary Markovian noises. Numerical validation through quantum state transfer and gate operations reveals near-unity fidelity ($99\%$) across diverse noise regimes, achieving orders-of-magnitude error suppression compared to target-only approaches. This framework bridges critical gaps between theoretical control design and experimental constraints, establishing a hardware-agnostic pathway toward fault-tolerant quantum technologies across platforms such as superconducting circuits, trapped ions, and solid-state qubits.
@@ I Introduction
Optimal control in open quantum systems must reconcile two competing objectives: driving the system toward a target state while simultaneously suppressing unwanted environmental couplings Weidner et al. (2025). To date, several control protocols—including dynamical decoupling Viola et al. (1999); Ryan et al. (2010), decoherence-free subspaces Lidar et al. (1998); Bacon et al. (2000), filter-function optimization Paz-Silva and Viola (2014); Cerfontaine et al. (2021), and quantum error correction Knill et al. (2000); Beale et al. (2018)—have demonstrated significant success in this endeavor. However, their effectiveness critically depends on prior knowledge of the mathematical structure of environmental noise Gordon et al. (2008); Schulte-Herbrüggen et al. (2011); Jing et al. (2013); Maloney et al. (2022); Liu et al. (2024); Chen et al. (2025), a requirement that is rarely satisfied in experimental settings due to the inherent difficulties of comprehensive noise characterization. This gap necessitates the development of universally applicable control strategies that remain effective without detailed knowledge of system-environment coupling.
@@
Here, we introduce a universally robust quantum control framework that substantially enhances operational fidelity in open quantum systems without requiring prior knowledge of noise characteristics. Our approach is built upon a key physical insight: external control drives not only manipulate the target system but also dynamically reconfigure the system-environment interaction itself Dann et al. (2018); Kallush et al. (2022); Di Meglio et al. (2024); Dann (2025). This dual effect is rigorously captured through a time-dependent master equation formalism, where the control fields instantaneously modulate the dissipative rates. Crucially, this description remains valid beyond adiabatic approximations and maintains thermodynamic consistency Kallush et al. (2022), establishing a physically grounded foundation for quantum control design. Through an analytical treatment of the master equation’s evolution operator, we derive an effective noise sensitivity metric that quantifies environmental susceptibility. This metric is then incorporated into a minimized objective functional defined exclusively by task-specific system observables, creating an inherently noise-agnostic optimization landscape. The resulting formulation provides provable robustness against arbitrary Markovian noises without requiring microscopic noise models—a mathematical manifestation of our protocol’s universality. We demonstrate the efficacy of our protocol through simulations of quantum state transfer and gate operations, optimized using the Chopped Random Basis (CRAB) algorithm Doria et al. (2011); Caneva et al. (2011); Müller et al. (2022). Both implementations achieve significant performance gains, obatining greater than 99% fidelity across a broad range of system-bath coupling strengths. Our framework’s platform-agnostic nature makes it broadly applicable to leading quantum hardware, including superconducting circuits Abdelhafez et al. (2020); Werninghaus et al. (2021), trapped ions Nebendahl et al. (2009); Shapira et al. (2018), and nitrogen-vacancy centers Poggiali et al. (2018); Vetter et al. (2024). This unified approach establishes a concrete pathway toward fault-tolerant quantum technologies, effectively bridging the divide between theoretical robustness and experimental feasibility.
@@ II.1 Model
where $\hat{H}_{\rm S}(t)=\hat{H}_{0}+\sum_{i}u_{i}(t)\hat{H}_{i}$ represents the noise-free system Hamiltonian, with $\hat{H}_{0}$ denoting the bare drift Hamiltonian, $\hat{H}_{i}$ the $i$-th control operator, and $u_{i}(t)$ the corresponding time-dependent control field. The bath Hamiltonian $\hat{H}_{\rm B}$ governs the environment’s free dynamics, which couples to the system via interaction Hamiltonian $\hat{H}_{\rm I}=g\sum_{\alpha}\hat{A}_{\alpha}\otimes\hat{B}_{\alpha}$ (the most general form), where $\hat{A}_{\alpha}$ and $\hat{B}_{\alpha}$ are Hermitian operators of the system and the bath, respectively, while $g$ characterizes their coupling strength. Under the Born-Markov approximation, tracing over the bath degrees of freedom yields the Gorini-Kossakowski-Lindblad-Sudarshan (GKLS) master equation Lindblad (1975); Gorini et al. (1976); Lindblad (1976) (in units where $\hbar=1$; see the Appendix for a detailed derivation)
@@
$\mathcal{D}[\hat{\rho}_{\rm S}]=\lambda\sum_{j}\kappa_{j}(t)\left[\hat{F}_{j}(t)\rho_{\rm S}(t)\hat{F}_{j}^{\dagger}(t)-\frac{1}{2}\left\{\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t),\rho_{\rm S}(t)\right\}\right]\,$, (3)
@@
where $\hat{U}_{\rm S}(t)$ is the system propagator evolving under ${\rm i}\partial_{t}\hat{U}_{\rm S}(t)=\hat{H}_{\rm S}(t)\hat{U}_{\rm S}(t)$ with initial condition $\hat{U}_{\rm S}(0)=\hat{\mathbb{I}}_{\rm S}$ (the identity operator for the system Hilbert space). The time-dependent decoherence rates $\kappa_{j}(t)$ read
@@
where $\eta_{j}^{\alpha}(t)=\left|\operatorname{\textnormal{Tr}}\left[{\hat{F}_{j}^{\dagger}(t)\hat{A}_{\alpha}}\right]\right|$ and $\gamma_{\alpha\alpha}\left[\omega_{j}(t)\right]$ is a function of the effective instantaneous Bohr frequencies $\omega_{j}(t)\left[={\rm d}\phi_{j}(t)/{\rm d}t\right]$.
Critically, the time-dependent master equation [Eq. (2)] explicitly accounts for how control drives in $\hat{H}_{\rm S}(t)$ modulate the dissipative dynamics. This results in control-dependent jump operators $\hat{F}_{j}(t)$ and time-varying decoherence rates $\kappa_{j}(t)$ Kallush et al. (2022). Consequently, the controller influences the system through two distinct channels: (i) Direct unitary steering via the coherent evolution ${\rm i}[\hat{H}_{\rm S}(t),\hat{\rho}_{\rm S}(t)]$ and (ii) Indirect dissipative control through the modulated dissipator $\mathcal{D}[\hat{\rho}_{\rm S}(t)]$. Remarkably, this framework enables the discovery of optimal controls $\hat{H}_{\rm S}(t)$ that simultaneously achieve target operations while actively suppressing environment-induced noise. As demonstrated in the subsequent section, such robust control protocols can be identified—even without detailed knowledge of system-bath couplings—through appropriate construction of the objective functional.
@@ II.2 Universally Robust Noise Mitigation
We now formulate a universal approach to suppress arbitrary Markovian noise by constructing a control-dependent cost function based on the time-dependent master equation [Eq. (2)]. This cost function depends solely on the designed control Hamiltonian $\hat{H}_{\rm S}(t)$ while remaining independent of specific noise channels.
For the sake of convenience, we first vectorize the master equation Eq. (2) by rewriting it in the Hilbert-Schmidt space Caves (1999); Andersson et al. (2007). This can be down by reshaping the density matrix $\hat{\rho}_{\rm S}(t)$ as a column vector, denoted as $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle$. It is straightforward to show that $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle$ satisfies the following Schrödinger-type equation:
@@
where the Lindbladian superoperator is vectorizd as
$\vec{\mathcal{L}}(t)=$ $-{\rm i}\left[\hat{H}_{\rm S}(t)\otimes\hat{\mathbb{I}}_{\rm S}-\hat{\mathbb{I}}_{\rm S}\otimes\hat{H}_{\rm S}(t)^{\mathrm{T}}\right]$ (7)
$+\lambda\sum_{j}\kappa_{j}(t)\left(\hat{F}_{j}(t)\otimes\hat{F}_{j}^{\ast}(t)-\frac{1}{2}\left[\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)\otimes\hat{\mathbb{I}}_{\rm S}+\hat{\mathbb{I}}_{\rm S}\otimes\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}^{\ast}(t)\right]\right)\,.$
The advantage of this vectorization technology is that it circumvents the complexity caused by the abstract superoperator and thus facilitates the perturbation analysis we applied subsequently.
By defining the shorthand $\mathbf{H}(t)=-{\rm i}[\hat{H}_{\rm S}(t)\otimes\hat{\mathbb{I}}_{\rm S}-\hat{\mathbb{I}}_{\rm S}\otimes\hat{H}^{\rm T}_{\rm S}(t)]$ and $\mathbf{F}^{j}(t)=\hat{F}_{j}(t)\otimes\hat{F}_{j}^{\ast}(t)-\frac{1}{2}[\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)\otimes\hat{\mathbb{I}}_{\rm S}+\hat{\mathbb{I}}_{\rm S}\otimes\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}^{\ast}(t)]$, the formal solution of Eq. (6) can be expressed as $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle=\mathbf{V}(t)|\hat{\rho}(0)\rangle\!\rangle$. Here,
@@
Robust quantum control requires insensitivity to first-order effects in the system-bath coupling strength $\lambda$. We therefore focus on the leading-order contribution, namely,
@@
The norm of Eq. (11) then quantifies the noise sensitivity, the minimization of which amounts to increasing the robustness of a control protocol. Aiming at acquiring a universal robustness, we first apply the Cauchy-Schwarz inequality to $\eta_{j}^{\alpha}(t_{1})$, yielding
@@
[…]
where $\mathbf{F}_{\alpha}=\int_{0}^{t}{\rm d}t_{1}\sum_{j}\gamma_{\alpha\alpha}\left[\omega_{j}(t_{1})\right]\widetilde{\mathbf{F}}^{j}(t_{1})$ and $\left\lVert\cdot\right\rVert$ denotes the Frobenius norm. In deriving the last inequality in Eq. (13), we have made use of the Cauchy-Schwarz inequality twice. Equation (13) suggests an effective noise sensitivity $D_{\rm eff}$, defined as
$D_{\rm eff}=\sqrt{\sum_{\alpha}\left\lVert\mathbf{F}_{\alpha}\right\rVert^{2}}\,.$ (14)
@@ II.3 Optimal Control Framework
We now demonstrate how our universal robustness measure $D_{\rm eff}$ integrates with quantum optimal control. The primary objective is to implement a target unitary operation $\hat{U}_{\rm tar}$ by optimizing the control fields $u_{i}(t)$ in the system Hamiltonian $\hat{H}_{\rm S}(t)$ [see details below Eq. (1)]. This is achieved by maximizing fidelity between $\hat{U}_{\rm tar}$ and the realized evolution $\hat{U}_{\rm S}(\tau)$, where $\tau$ is total evolution time. We distinguish two control paradigms: (i) For the task of pure state transfer from an initial state $\hat{\rho}_{\rm i}$ to a target state $\hat{\rho}_{\rm tar}$, such fidelity reduces to
@@
where $\hat{\rho}_{\rm f}^{n}=\hat{U}_{\rm S}(\tau)\hat{\rho}_{\rm i}^{n}\hat{U}^{\dagger}_{\rm S}(\tau)$, $\hat{\rho}_{\rm tar}^{n}=\hat{U}_{\rm tar}\hat{\rho}_{\rm i}^{n}\hat{U}^{{\dagger}}_{\rm tar}$, and $N$ is the dimension of system’s Hilbert space. Note that we have excluded the identity since it is preserved in the CPTP map and this definition guarantees $\mathcal{F}_{\text{gate}}\leq 1$. In the following, we optimize the infidelity $\mathcal{J}_{0}=1-\mathcal{F}_{\rm state/gate}$ for numerical efficiency, converting fidelity maximization to minimization of a positive definite functional.
Robust optimal control additionally requires that the control process owns immunity to environment noise. We then face a multi-objective optimization task in which the total objective functional to be minimized can be organized as
@@
We implement both control strategies using the CRAB algorithm Doria et al. (2011); Caneva et al. (2011); Müller et al. (2022), which offers efficient optimization with experimentally advantageous features like inherent pulse smoothness and a reduced parameter space. The specific parameterization is $u_{i}(t)=\exp(-[(t-\tau/2)/(2\sigma)]^{2})\sum_{k=1}^{M}c_{k}\sin(\nu_{k}t)$, where $c_{k}$} are optimization coefficients, $\sigma$ controls the Gaussian envelope width, and $\nu_{k}=k\pi/\tau$ ($k=1,\dots,M$) are fixed frequencies (in the following, we choose $\sigma=\tau/4$ and $M=10$). We optimize $c_{k}$} using a quasi-Newton algorithm to minimize $\mathcal{J}$.
Figure 1: Universally robust quantum state transfer in a two-level system. Quantum state transfer from initial state $\hat{\rho}_{-x}$ [Bloch vector (1,0,0)(-1,0,0)] to target state $\hat{\rho}_{+x}$ [Bloch vector $(+1,0,0)$] under environmental noise. Gray dash-dotted lines: target-only control ($c=0$). Orange solid lines: universally robust control ($c=10^{-2}$). (a) State fidelity $\mathcal{F}_{\rm state}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{z}$). Inset: Log-log infidelity ($1-\mathcal{F}_{\rm state}$). (b) $\mathcal{F}_{\rm state}$ versus $\lambda$ for generic noise ($\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$; $\mathbf{n}$: random unit vector). Data averaged over 20 realizations. Inset: Corresponding infidelity. (c) Optimized control fields $u_{x}(t)$ and $u_{y}(t)$ for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. (d) State evolution trajectories on Bloch sphere for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. Initial state: black square. Final states: blue dot (target-only) and green diamond (universally robust control). Parameters: Energy splitting $\Delta=3\times 10^{-3}$ a.u., inverse temperature $\beta=1/\Delta$.

### v2

@@ Abstract
Mitigating noise-induced decoherence is the central challenge in controlling open quantum systems. While existing robust protocols often require precise noise models, we introduce a universal framework for noise-agnostic quantum control that achieves high-fidelity operations without prior environmental noise characterization. This framework capitalizes on the dynamical modification of the system-environment coupling through control drives, an effect rigorously encoded in the dynamical equation. Since the derived noise sensitivity metric remains independent of the coupling details between the system and the environment, our protocol demonstrates robustness against arbitrary Markovian noises within the first-order weak coupling approximation. Numerical validation through quantum state transfer and gate operations reveals near-unity fidelity across diverse noise regimes, achieving orders-of-magnitude error suppression compared to target-only approaches. This framework bridges critical gaps between theoretical control design and experimental constraints, establishing a broadly applicable strategy for high-fidelity quantum information processing across platforms such as superconducting circuits, trapped ions, and solid-state qubits.
@@ I Introduction
Optimal control in open quantum systems must reconcile two competing objectives: driving the system toward a target state while simultaneously suppressing unwanted environmental couplings Weidner et al. (2025). To date, several control protocols—including dynamical decoupling Viola et al. (1999); Ryan et al. (2010), decoherence-free subspaces Lidar et al. (1998); Bacon et al. (2000), and filter-function optimization Paz-Silva and Viola (2014); Cerfontaine et al. (2021)—have demonstrated significant success in this endeavor. However, their effectiveness critically depends on prior knowledge of the mathematical structure of environmental noise Gordon et al. (2008); Schulte-Herbrüggen et al. (2011); Jing et al. (2013); Maloney et al. (2022); Liu et al. (2024); Chen et al. (2025), a requirement that is rarely satisfied in experimental settings due to the inherent difficulties of comprehensive noise characterization. This gap necessitates the development of universally applicable control strategies that remain effective without detailed knowledge of system-environment coupling.
@@
Here, we introduce a universally robust quantum control framework that substantially enhances operational fidelity in open quantum systems without requiring prior knowledge of noise characteristics. Our approach is built upon a key physical insight: external control drives not only manipulate the target system but also dynamically reconfigure the system-environment interaction itself Dann et al. (2018); Kallush et al. (2022); Di Meglio et al. (2024); Dann (2025). This dual effect is rigorously captured through a time-dependent master equation formalism, where the control fields instantaneously modulate the dissipative rates. Crucially, this description remains valid beyond adiabatic approximations and maintains thermodynamic consistency Kallush et al. (2022), establishing a physically grounded foundation for quantum control design. Through an analytical treatment of the master equation’s evolution operator, we derive an effective noise sensitivity metric that quantifies environmental susceptibility. This metric is then incorporated into a minimized objective functional defined exclusively by task-specific system observables, creating an inherently noise-agnostic optimization landscape. The resulting formulation provides robustness against arbitrary Markovian noises within the first-order weak coupling approximation without requiring microscopic noise models—a mathematical manifestation of our protocol’s universality. We demonstrate the efficacy of our protocol through simulations of quantum state transfer and gate operations, optimized using the Chopped Random Basis (CRAB) algorithm Doria et al. (2011); Caneva et al. (2011); Müller et al. (2022). Both implementations achieve significant performance gains, obtaining greater than 98% fidelity across a broad range of system-bath coupling strengths. Our framework’s versatile nature makes it broadly applicable to leading quantum hardware, including superconducting circuits Abdelhafez et al. (2020); Werninghaus et al. (2021), trapped ions Nebendahl et al. (2009); Shapira et al. (2018), and nitrogen-vacancy centers Poggiali et al. (2018); Vetter et al. (2024). This unified approach establishes a concrete approach for robust quantum control, effectively bridging the divide between theoretical robustness and experimental feasibility.
@@ II.1 Model
where $\hat{H}_{\rm S}(t)=\hat{H}_{0}+\sum_{i}u_{i}(t)\hat{H}_{i}$ represents the noise-free system Hamiltonian, with $\hat{H}_{0}$ denoting the bare drift Hamiltonian, $\hat{H}_{i}$ the $i$-th control operator, and $u_{i}(t)$ the corresponding time-dependent control field. The bath Hamiltonian $\hat{H}_{\rm B}$ governs the environment’s free dynamics, which couples to the system via interaction Hamiltonian $\hat{H}_{\rm I}=g\sum_{\alpha}\hat{A}_{\alpha}\otimes\hat{B}_{\alpha}$ (the most general form), where $\hat{A}_{\alpha}$ and $\hat{B}_{\alpha}$ are Hermitian operators of the system and the bath, respectively, while $g$ characterizes their coupling strength. Under the Born-Markov approximation, tracing over the bath degrees of freedom yields the Gorini-Kossakowski-Lindblad-Sudarshan (GKLS) master equation Lindblad (1975); Gorini et al. (1976); Lindblad (1976); Pearle (2012); Manzano (2020) (in units where $\hbar=1$)
@@
$\mathcal{D}[\hat{\rho}_{\rm S}(t)]=\lambda\sum_{j}\kappa_{j}(t)\left[\hat{F}_{j}(t)\hat{\rho}_{\rm S}(t)\hat{F}_{j}^{\dagger}(t)-\frac{1}{2}\{\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t),\hat{\rho}_{\rm S}(t)\}\right]\,$, (3)
@@
[…]
@@
where $\eta_{j}^{\alpha}(t)=\left|\operatorname{\textnormal{Tr}}\left[{\hat{F}_{j}^{\dagger}(t)\hat{A}_{\alpha}}\right]\right|$ and $\gamma_{\alpha\alpha}[\omega_{j}(t)]=\int_{-\infty}^{\infty}{\rm d}s\,{\rm e}^{-{\rm i}\omega_{j}(t)s}\,{\rm Tr}_{\rm B}\left[\tilde{B}_{\alpha}(s)\tilde{B}_{\alpha}(0)\hat{\rho}_{\rm B}\right]$ with $\hat{\rho}_{\rm B}$ the density operator of the bath (see Method A for a detailed discussion on $\gamma_{\alpha\alpha}[\omega_{j}(t)]$). Here, $\tilde{B}_{\alpha}(s)=\hat{U}_{\rm B}^{\dagger}(s)\hat{B}_{\alpha}\hat{U}_{B}(s)$ are the bath operators in the interaction picture and $\omega_{j}(t)\left[={\rm d}\phi_{j}(t)/{\rm d}t\right]$ account for the effective instantaneous Bohr frequencies.
Critically, the time-dependent master equation explicitly accounts for how control drives in $\hat{H}_{\rm S}(t)$ modulate the dissipative dynamics. This results in control-dependent jump operators $\hat{F}_{j}(t)$ and time-varying decoherence rates $\kappa_{j}(t)$ Kallush et al. (2022). Consequently, the controller influences the system through two distinct channels: (i) Direct unitary steering via the coherent evolution ${\rm i}[\hat{H}_{\rm S}(t),\hat{\rho}_{\rm S}(t)]$ and (ii) Indirect dissipative control through the modulated dissipator $\mathcal{D}[\hat{\rho}_{\rm S}(t)]$. Remarkably, this framework enables the discovery of optimal controls $\hat{H}_{\rm S}(t)$ that simultaneously achieve target operations while actively suppressing environment-induced noise. As demonstrated in the subsequent section, such robust control protocols can be identified—even without detailed knowledge of system-bath couplings—through appropriate construction of the objective functional.
@@ II.2 Universally Robust Noise Mitigation
We now formulate a universal approach to suppress arbitrary Markovian noise by constructing a control-dependent cost function based on the time-dependent master equation. This cost function depends solely on the designed control Hamiltonian $\hat{H}_{\rm S}(t)$ while remaining independent of specific noise channels.
For the sake of convenience, we first vectorize the master equation Eq. (2) by rewriting it in the Hilbert-Schmidt space Caves (1999); Andersson et al. (2007). This can be done by reshaping the density matrix $\hat{\rho}_{\rm S}(t)$ as a column vector, denoted as $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle$. It is straightforward to show that $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle$ satisfies the following Schrödinger-type equation:
@@
where the Lindbladian superoperator is vectorized as
$\vec{\mathcal{L}}(t)=$ $-{\rm i}\left[\hat{H}_{\rm S}(t)\otimes\hat{\mathbb{I}}-\hat{\mathbb{I}}\otimes\hat{H}_{\rm S}^{\rm T}(t)\right]$ (7)
$+\lambda\sum_{j}\kappa_{j}(t)\left(\hat{F}_{j}(t)\otimes\hat{F}_{j}^{\ast}(t)-\frac{1}{2}\left[\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)\otimes\hat{\mathbb{I}}+\hat{\mathbb{I}}\otimes(\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t))^{\rm T}\right]\right)\,.$
This Hilbert-Schmidt vectorization is mathematically equivalent to the Bloch representation but is employed here as it transforms abstract superoperators into concrete $N^{2}\times N^{2}$ matrices, which is computationally convenient for perturbation analysis and numerical optimization we applied subsequently.
By defining the shorthand $\mathbf{H}(t)=-{\rm i}[\hat{H}_{\rm S}(t)\otimes\hat{\mathbb{I}}-\hat{\mathbb{I}}\otimes\hat{H}_{\rm S}^{\rm T}(t)]$ and $\mathbf{F}^{j}(t)=\hat{F}_{j}(t)\otimes\hat{F}_{j}^{\ast}(t)-\frac{1}{2}[\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)\otimes\hat{\mathbb{I}}+\hat{\mathbb{I}}\otimes(\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t))^{\rm T}]$, the formal solution of Eq. (6) can be expressed as $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle=\mathbf{V}(t)|\hat{\rho}(0)\rangle\!\rangle$. Here,
@@
Robust quantum control requires insensitivity to first-order effects in the system-bath coupling strength $\lambda$ Not . We therefore focus on the leading-order contribution, namely,
@@
The norm of Eq. (11) then quantifies the noise sensitivity, the minimization of which amounts to increasing the robustness of a control protocol. Aiming at acquiring a universal robustness, we first apply the Cauchy-Schwarz inequality to $\eta_{j}^{\alpha}(t_{1})]^{2}$, yielding
@@
$\left\lVert\left.\frac{{\rm d}\mathbf{U}_{\mathrm{err}}(t)}{{\rm d}\lambda}\right|_{\lambda=0}\right\rVert\leq\sum_{\alpha}\operatorname{\textnormal{Tr}}\left[{\hat{A}_{\alpha}^{\dagger}\hat{A}_{\alpha}}\right]\cdot\left\lVert\mathbf{F}_{\alpha}\right\rVert\leq\sqrt{\sum_{\alpha}\operatorname{Tr}^{2}\left[\hat{A}_{\alpha}^{\dagger}\hat{A}_{\alpha}\right]}\cdot\sqrt{\sum_{\alpha}\left\lVert\mathbf{F}_{\alpha}\right\rVert^{2}}\,$, (13)
where $\mathbf{F}_{\alpha}=\int_{0}^{t}{\rm d}t_{1}\sum_{j}\gamma_{\alpha\alpha}\left[\omega_{j}(t_{1})\right]\widetilde{\mathbf{F}}^{j}(t_{1})$ and $\left\lVert\mathbf{F}_{\alpha}\right\rVert=\sqrt{\text{Tr}[\mathbf{F}^{\dagger}_{\alpha}\mathbf{F}_{\alpha}]}$ denotes the Frobenius norm. In deriving the last inequality in Eq. (13), we have made use of the Cauchy-Schwarz inequality twice. Equation (13) suggests an effective noise sensitivity $D_{\rm eff}$, defined as (see Method B for a detailed recipe to obtain $D_{\rm eff}$)
$D_{\mathrm{eff}}=\sqrt{\sum_{\alpha}\left\lVert\mathbf{F}_{\alpha}\right\rVert^{2}}\,.$ (14)
@@ II.3 Optimal Control Framework
We now demonstrate how our universal robustness measure $D_{\rm eff}$ integrates with quantum optimal control. The primary objective is to implement a target unitary operation $\hat{U}_{\rm tar}$ by optimizing the control fields $u_{i}(t)$ in the system Hamiltonian $\hat{H}_{\rm S}(t)$ [see details below Eq. (1)]. This is achieved by maximizing fidelity between $\hat{U}_{\rm tar}$ and the realized evolution $\hat{U}_{\rm S}(\tau)$, where $\tau$ is total evolution time. We distinguish two control paradigms: (i) For the task of state transfer from an initial pure state $\hat{\rho}_{\rm i}$ to a target pure state $\hat{\rho}_{\rm tar}$, such fidelity reduces to
@@
where $\hat{\rho}_{\rm f}^{n}=\hat{U}_{\rm S}(\tau)\hat{\rho}_{\rm i}^{n}\hat{U}^{\dagger}_{\rm S}(\tau)$, $\hat{\rho}_{\rm tar}^{n}=\hat{U}_{\rm tar}\hat{\rho}_{\rm i}^{n}\hat{U}^{{\dagger}}_{\rm tar}$. Note that we have excluded the identity since it is preserved in the CPTP map and this definition guarantees $\mathcal{F}_{\text{gate}}\leq 1$. In the following, we optimize the infidelity $\mathcal{J}_{0}=1-\mathcal{F}_{\rm state/gate}$ for numerical efficiency, converting fidelity maximization to minimization of a positive definite functional.
Robust optimal control additionally requires that the control process is immune environment noise. We then face a multi-objective optimization task in which the total objective functional to be minimized can be organized as Poggi et al. (2024)
@@
We implement both control strategies using the CRAB algorithm Doria et al. (2011); Caneva et al. (2011); Müller et al. (2022), which offers efficient optimization with experimentally advantageous features like inherent pulse smoothness and a reduced parameter space. The specific parameterization is $u_{i}(t)=\exp(-[(t-\tau/2)/(2\sigma)]^{2})\sum_{k=1}^{M}c_{k}\sin(\nu_{k}t)$, where $c_{k}$} are optimization coefficients, $\sigma$ controls the Gaussian envelope width, $\tau$ denotes the total control time, and $\nu_{k}=k\pi/\tau$ ($k=1,\dots,M$) are fixed frequencies (in the following, we choose $\sigma=\tau/4$ and $M=10$). We optimize $c_{k}$} using a quasi-Newton algorithm to minimize $\mathcal{J}$.